OpenStudy (anonymous):

Find the sum and difference functions f+g and f−g for the functions given: a. f(x)=x2−2x−2 and g(x)=3x+2 b. f(x)=9−x2 and g(x)=5x2+2x+1 c. f(x)=2x square root 5+4 and g(x)=x2+2x square root 5−1

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

step by step answer please :)

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x)=x^2-2x-2+3x+2=x^2+x

OpenStudy (anonymous):

is that everything ?

OpenStudy (anonymous):

yes, similarly for others.

OpenStudy (anonymous):

can you do part b please ?

OpenStudy (anonymous):

b. i solve f-g (f-g)(x)=f(x)-g(x)=9-x^2-(5x^2+2x+1)=9-x^2-5x^2-2x-1=-6x^2-2x+8

OpenStudy (anonymous):

6x^2-2x+8 is the final answer right?

OpenStudy (anonymous):

no ,-6x^2-2x+8

OpenStudy (anonymous):

oh yes sorry

OpenStudy (anonymous):

and how about part c? Im doing part e,f,g on my own dw

OpenStudy (anonymous):

the square roots confuse me @surjithayer

OpenStudy (anonymous):

ok can you keep going with it ??

OpenStudy (anonymous):

f(x)=2x(sqrt5)+4,g(x)=x^2+2x(sqrt5)-1 (f+g)(x)=f(x)+g(x)=2x(sqrt5)+4+x^2+2x(sqrt5)-1=x^2+4x(sqrt5)+3

OpenStudy (anonymous):

thanks so much !! :D

OpenStudy (anonymous):

yw